OpenStudy (anonymous):

Write a system of two first degree equations involving the variables x and y that represent the conditions stated in the problem. Solve and graph The sum of two numbers is 8. The difference of these numbers is 2. Find the numbers

OpenStudy (anonymous):

Can you write an equation using "the sum of two numbers is 8"

OpenStudy (anonymous):

No i am having alot of trouble

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

Let's start with an example. The sum of two number is 4: x + y = 4 The difference of two number sis 2: x - y = 2 Find the numbers x + y = 4 --> x = 4 - y x - y = 2 --> substitute x = 4 - y in 4 - y - y = 2 4 - 2y = 2 4 - 2 = 2y 2 = 2y 1 = y Substitute this back into any of the original equations x + y = 4 x + 1 = 4 x = 3 Follow this method for your current problem.

OpenStudy (anonymous):

Ok i will try wait one minute

OpenStudy (anonymous):

Y =3 and x=5

OpenStudy (anonymous):

Perfect

OpenStudy (anonymous):

Thank you